OpenStudy (anonymous):

what is the derivative of 2(512/x)?

OpenStudy (anonymous):

\[\LARGE y=2\left( \frac{512}{x}\right)\] \[\LARGE y=2\times 512\left( \frac{1}{x}\right)\] \[\LARGE y=1024 \left( x^{-1 }\right)\] \[\LARGE y'=1024 \left( x^{-1 }\right)'\] \[\LARGE y'=1024 \times(-1)\times \left( x^{-1-1 }\right)\] \[\LARGE y'=-1024 \times \left( x^{-2 }\right)\] \[\LARGE y'=-1024 \times \left( \frac{1}{x^{2 }}\right)\]

OpenStudy (anonymous):

thank you. i always get confused when the variable is in the denominator

OpenStudy (anonymous):

You're welcome... just move numbers aside. And the rest is easy :) ...